OpenStudy (anonymous):

Convert each expression into exponential form. 1. 1/x^2 2. 2/x^3 3. 4/x^2 4.1/2√(x) 5. 3/ 2√(x^3)

OpenStudy (ybarrap):

$$ 1. \frac 1{x^2}=x^{-2} \\ 2. \frac 2{x^3} =2x^{-3}\\ 3. \frac 4{x^2} =4x^{-2}\\ 4. \frac 1{2\sqrt x} =\frac {1}{2x^{1/2}}=\frac {x^{-1/2}}{2}\\\\ 5. \frac 3{ 2\sqrt{x^3}}=\frac {3}{2x^{3/2}}=\frac {3x^{-3/2}}{2}\\\\ $$ Questions?

OpenStudy (anonymous):

oh nice, awesome. nope! thanks man

OpenStudy (anonymous):

oh also what is √3(x^2)?

OpenStudy (ybarrap):

$$ \huge \sqrt{3x^2}=\left ( 3x^2\right )^{\frac 1 2}= 3^{\frac 1 2}x^{\frac 2 2}=3^{\frac 1 2}x^1=3^{\frac 1 2}x\\\huge=\sqrt 3x=x\sqrt 3 $$